OpenStudy (anonymous):

Please! help A die is rolled. Find each probability. 1). P(multiple of 3 or 4) 2). P(multiple of 2 or 3)

OpenStudy (anonymous):

How many multiples of 3 are there on a single die? How many multiples of 4?

OpenStudy (anonymous):

1. multiple of 3 is 3,6 and multiple of 4 is 4 so numbers are 3,4,6 p(multiple of 3 or 4) = 3/6 =1/2

OpenStudy (anonymous):

2. multiple of 2 are 2,4,6 and multiple of 3 are 3,6 so, common numbers are 2,3,4,6 P(multiple of 2 or 3) =4/6 =2/3

OpenStudy (anonymous):

3*2=6 4*1=4?

OpenStudy (anonymous):

3 is also a multiple of 3

OpenStudy (anonymous):

3*1=3?

OpenStudy (anonymous):

there are only two number multiple of 3 that is 3 and 6

OpenStudy (anonymous):

and there is only one multiple of 4 in a dice that is 4 itself

OpenStudy (anonymous):

than we have 3,6,4 ?

OpenStudy (anonymous):

yes , only number multiple for multiple of 3 or 4

OpenStudy (anonymous):

are 3,4,6

OpenStudy (anonymous):

so,p(multiple of 3 or 4) = 3/6 =1/2

OpenStudy (anonymous):

how about 2)?

OpenStudy (anonymous):

2. multiple of 2 are 2,4,6 and multiple of 3 are 3,6

OpenStudy (anonymous):

so, common numbers are 2,3,4,6

OpenStudy (anonymous):

so, P(multiple of 2 or 3) =4/6 =2/3

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

my pleasure:)